A rectangular swimming pool is 50 ft. wide, 100 ft. long, and 10 ft. deep. The pool will be filled to a line 2 ft. from the top. The water company charges $2.25/1000 ft3. How much does it cost to fill the pool?
Answer by checkley77(12844)

50*100*8=40,000 ft^3 is the volume of water in the pool.
(40,000/1,000)2.25=40*2,25=$90 price to fill the pool.